A large group of students took a test in physics and have a mean of 70 and a standard deviation of 10.If we approximate the distribution of these grades by a normal distribution, what percentage of the students


a) scored higher than 80?


b) should pass the test(grades>=60)?


c) should fail the test(grades<60)?

1
Expert's answer
2022-03-07T17:08:01-0500

Let "X" be the random variable denoting the mark in the test scored by the students.

Given, "\mu = 70, \sigma = 10." Let "z = \dfrac{X - \mu}{\sigma} = \dfrac{X - 70}{10}"


a. "\begin{aligned}
P(\text{students score more than 80)} &= P(X > 80)\\
&= P\left(\dfrac{X-\mu}{\sigma} > \dfrac{80 - 70}{10}\right)\\
&= P(z > 1)\\
&= 1 - P(z < 1)\\ 
& = 1 - 0.8413\quad(\text{Using Normal distribution table})\\
&= 0.1587
\end{aligned}"

Therefore, 15.87% of students scored above 80 marks.


b.

"\begin{aligned}
P(\text{students pass the test)} &= P(X \ge 60)\\
&= P\left(\dfrac{X-\mu}{\sigma} \ge \dfrac{60 - 70}{10}\right)\\
&= P(z \ge -1)\\
&= 1 - P(z < -1)\\ 
& = 1 - 0.1587\quad(\text{Using Normal distribution table})\\
&= 0.8413
\end{aligned}"

Therefore, 84.13% of students passed in the test.


c.

"\begin{aligned}
P(\text{students fail the test)} &= P(X < 60)\\
&= P\left(\dfrac{X-\mu}{\sigma} < \dfrac{60 - 70}{10}\right)\\
&= P(z < -1)\\
& = 0.1587\quad(\text{Using Normal distribution table})\\
\end{aligned}"

Therefore, 15.87% of students failed in the test.